# Enhancing Biomedical Knowledge Discovery for Diseases: An End-To-End Open-Source Framework

### v1

@@ 6 Related Work
Information Extraction Datasets. Several biomedical datasets aim to enhance Information Extraction (IE) system development (Huang et al., 2024; Detroja et al., 2023; Nasar et al., 2021; Theodoropoulos et al., 2021), typically focusing on one or a few entity types and their interactions. AIMed (Bunescu et al., 2005), BioInfer (Pyysalo et al., 2007), and BioCreative II PPI IPS (Krallinger et al., 2008) formulate protein-protein interactions. The chemical-protein and chemical-disease interactions are modeled by DrugProt (Miranda et al., 2021) and BC5CDR (Li et al., 2016), respectively. ADE (Gurulingappa et al., 2012), DDI13 (Herrero-Zazo et al., 2013), and n2c2 2018 ADE (Henry et al., 2020) include drug-ADE (adverse drug effect) and drug-drug interactions. EMU (Doughty et al., 2011), GAD (Bravo et al., 2015) and RENET2 (Su et al., 2021) contain relations between genes and diseases. N-ary (Peng et al., 2017) incorporates drug-gene mutation interactions. The task of event extraction is illustrated by GE09 (Kim et al., 2009), GE11 (Kim et al., 2011), and CG (Pyysalo et al., 2013). DDAE (Lai et al., 2019) includes disease-disease associations. BioRED (Luo et al., 2022) focuses on document-level relations for various entities. Unlike these datasets, ReDReS and ReDAD focus on RS and AD, include entities of up to 82 different semantic types, and model the semantic relation between them.
@@
Knowledge Discovery. Gottlieb et al. (2011) present PREDICT, a method for ranking potential drug-disease associations to predict drug indications. Romano et al. (2024) release AlzKB, a heterogeneous graph knowledge base for AD, constructed using external data sources and describing various medical entities (e.g., chemicals, genes). Other graph-based efforts model knowledge around AD for tasks such as drug repurposing (Hsieh et al., 2023; Daluwatumulle et al., 2022; Nian et al., 2022), gene identification (Binder et al., 2022), or as general knowledge repositories (Sügis et al., 2019). Another paradigm for knowledge discovery is the open information extraction (OIE) setup (Mausam et al., 2012; Etzioni et al., 2008), which faces challenges such as data consistency, performance evaluation, and semantic drift (Zhou et al., 2022). Research efforts (Wang et al., 2018; de Silva et al., 2017; Nebot and Berlanga, 2014; Movshovitz-Attias and Cohen, 2012; Nebot and Berlanga, 2011) aim to address these issues and extract knowledge with little or no supervision. Advances in literature-based discovery (Gopalakrishnan et al., 2019; Thilakaratne et al., 2019) try to identify novel medical entity relations using graph-based (Kilicoglu et al., 2020; Nicholson and Greene, 2020), machine learning (Zhao et al., 2021; Lardos et al., 2022), and co-occurrence methods (Kuusisto et al., 2020; Millikin et al., 2023). Tian et al. (2024) stress the potential of large language models (LLMs) to summarize, simplify, and synthesize medical evidence (Peng et al., 2023; Tang et al., 2023; Shaib et al., 2023), suggesting that LLMs may have encoded biomedical knowledge (Singhal et al., 2023). To exploit this potential, we explore constructing LM representations for knowledge discovery. To the best of our knowledge, no systematic approach assembles knowledge about RS. Unlike previous work, we introduce a, in principle, disease-agnostic framework, to acquire knowledge about RS and AD starting from raw text.
@@ Appendix A Data Pipeline: Additional Information
To facilitate effective abstract retrieval, we implement an iterative approach to circumvent the API’s limitation of retrieving only 10,000 article IDs per query. This iterative process enables us to access a comprehensive set of PubMed IDs (PMIDs) related to the query. The detailed list of the 82 semantic types of the MetaMapLite-based pipeline is presented in Table 4. In addition to the MetaMapLite-based pipeline, we propose a second pipeline that is based on ScispaCy (Apache License 2.0) (Fig. 5). The difference lies in the selection of entity extractors and linkers that map the extracted entities to knowledge schemes. Unlike MetaMapLite, which adopts an integrated approach where mention extraction and linking are performed simultaneously in a single step and focuses on UMLS mapping, allowing for more precise and targeted extraction of entities, ScispaCy serves a broader range of Natural Language Processing (NLP) tasks. After the retrieval of the abstracts that is described in subsection 2.1, the following steps are executed: Knowledge schema and linker generation, Mention extraction, Entity linking. and Sampling of linked identifiers.
@@
Entity linking. This process enhances the semantic understanding of the extracted entities, facilitates standardization, which is a key issue in the biomedical field (Bettencourt-Silva et al., 2012; Theodoropoulos et al., 2023), and promotes interoperability with external resources by associating the entities with specific concepts in supported knowledge schemes. Each entity is subjected to a linking process where we attempt to map it to concepts within supported knowledge bases or vocabularies. If a match is found, the entity is assigned a unique identifier, referred to as a CUI, corresponding to the specific concept in the knowledge schema. As entities may be linked to multiple knowledge sources, we merge the extracted CUIs obtained from the different linkers. This consolidation process ensures that each entity is associated with a comprehensive set of identifiers, encompassing diverse perspectives and representations across various knowledge schemes.
@@
$sam\_sent\_2\leftarrow sample(\mathcal{IP},n/2)$
Return $sam\_sent\_1$, $sam\_sent\_2$
@@ Appendix D Localized Context Vector
Appendix E Baseline Performance: Lower Bound
@@
Appendix F Distantly Supervised Datasets
@@
  • LaMReDA (PubMedBERT large) with the relation representation $R_{A}$ (Eq. 9) for the RS corpus.
  • LaMReDA (PubMedBERT large) with the relation representation $R_{J}$ (Eq. 18) for the AD corpus.
@@
  • LaMReDA (PubMedBERT large) with the relation representation $R_{F}$ (Eq. 14) for the RS corpus.
  • LaMReDA (PubMedBERT large) with the relation representation $R_{F}$ (Eq. 14) for the AD corpus.
@@
Weakly Supervised Setup. The task formulation remains the same as described in section 3 of the paper. The train sets of ReDReS and ReDAD are replaced by DiSReDReS and DiSReDAD, respectively. The development and test sets remain the same (Tab. 1). To provide a benchmark, we train the LaMReDA (PubMedBERT base) with the relation representation $R_{A}$ (Eq. 9) for 10 epochs utilizing the ADAM optimizer with learning rate 10-5. The batch is set to 32. The experiments are repeated 10 times with different seeds and the best scores are retained based on the performance on the development set.
@@
In the supervised setup (Tab. 1), LaMReDA (PubMedBERT base) with the $R_{A}$ representation achieves 90.72% and 88.31% F1-score in the binary setup on ReDReS and ReDAD, respectively (Tab. 3). Multi-class macro F1-scores range from 74.52% (micro: 74.49%) to 77.34% (micro: 77.64%) for ReDReS and ReDAD, accordingly (Tab. 3). Table 5 presents the benchmark performance in the binary and multi-class setup for both datasets. Notably, the performance is improved in the weakly supervised setup, indicating the robustness of LaMReDA, when trained with noisy data, and highlighting the quality of the silver labels of DiSReDReS and DiSReDAD.
Appendix G Probing: Additional Experiments
We use the same experimental setup as described in subsection 3.3 and the experiments are conducted in the 5-fold cross-validation setting. To provide an inclusive probing analysis on ReDReS, we incorporate additional probing results in this section. Figures 8 and 9 present the experiments in the multi-class setup using PubMedBERT base. Additionally, aiming to explore the probing capabilities of PubMedBERT large, we include the results of further experiments in Figures 10, 11, and 12. These experiments investigate the model’s performance in detecting semantic relations, comparing the representations and attention mechanisms at different layers and heads to understand how well the larger LM can discern complex relationships in the biomedical text.
@@ Appendix G Probing: Additional Experiments
Figure 8: ReDReS Probing (Multi-class setup, Macro evaluation) (PubMedBERT base): Examines LaMReDA/LaMReDM relation representations ($R_{D}$, $R_{O}$, $R_{P}$) and attention scores from each layer and explores average attention scores of tokens corresponding to each entity towards the other entity across attention heads. Top boundary: best LaMReDA and LaMReDM performance (Tab. 3). Second boundary: classifier with average attention scores across all layers as input.
@@
Figure 9: ReDReS Probing (Multi-class setup, Micro evaluation) (PubMedBERT base): Examines LaMReDA/LaMReDM relation representations ($R_{D}$, $R_{O}$, $R_{P}$) and attention scores from each layer and explores average attention scores of tokens corresponding to each entity towards the other entity across attention heads. Top boundary: best LaMReDA and LaMReDM performance (Tab. 3). Second boundary: classifier with average attention scores across all layers as input.
@@
Figure 10: ReDReS Probing (Binary setup) (PubMedBERT large): Examines LaMReDA/LaMReDM relation representations ($R_{D}$, $R_{O}$, $R_{P}$) and attention scores from each layer and explores average attention scores of tokens corresponding to each entity towards the other entity across attention heads. Top boundary: best LaMReDA and LaMReDM performance (Tab. 3). Second boundary: classifier with average attention scores across all layers as input.
@@
Figure 11: ReDReS Probing (Multi-class setup, Macro evaluation) (PubMedBERT large): Examines LaMReDA/LaMReDM relation representations ($R_{D}$, $R_{O}$, $R_{P}$) and attention scores from each layer and explores average attention scores of tokens corresponding to each entity towards the other entity across attention heads. Top boundary: best LaMReDA and LaMReDM performance (Tab. 3). Second boundary: classifier with average attention scores across all layers as input.
@@
Figure 12: ReDReS Probing (Multi-class setup, Micro evaluation) (PubMedBERT large): Examines LaMReDA/LaMReDM relation representations ($R_{D}$, $R_{O}$, $R_{P}$) and attention scores from each layer and explores average attention scores of tokens corresponding to each entity towards the other entity across attention heads. Top boundary: best LaMReDA and LaMReDM performance (Tab. 3). Second boundary: classifier with average attention scores across all layers as input.
Appendix H Detailed Annotation Guidelines

### v2

@@ 6 Related Work
Information Extraction Datasets. Several biomedical datasets aim to enhance Information Extraction (IE) system development (Theodoropoulos et al., 2021; Nasar et al., 2021; Detroja et al., 2023; Huang et al., 2024), typically focusing on one or a few entity types and their interactions. AIMed (Bunescu et al., 2005), BioInfer (Pyysalo et al., 2007), and BioCreative II PPI IPS (Krallinger et al., 2008) formulate protein-protein interactions. The chemical-protein and chemical-disease interactions are modeled by DrugProt (Miranda et al., 2021) and BC5CDR (Li et al., 2016), respectively. ADE (Gurulingappa et al., 2012), DDI13 (Herrero-Zazo et al., 2013), and n2c2 2018 ADE (Henry et al., 2020) include drug-ADE (adverse drug effect) and drug-drug interactions. EMU (Doughty et al., 2011), GAD (Bravo et al., 2015) and RENET2 (Su et al., 2021) contain relations between genes and diseases. N-ary (Peng et al., 2017) incorporates drug-gene mutation interactions. The task of event extraction is illustrated by GE09 (Kim et al., 2009), GE11 (Kim et al., 2011), and CG (Pyysalo et al., 2013). DDAE (Lai et al., 2019) includes disease-disease associations. BioRED (Luo et al., 2022a) focuses on document-level relations for various entities. Unlike these datasets, ReDReS and ReDAD focus on RS and AD, include entities of up to 82 different semantic types, and model the semantic relation between them.
@@
Knowledge Discovery. Gottlieb et al. (2011) present PREDICT, a method for ranking potential drug-disease associations to predict drug indications. Romano et al. (2024) release AlzKB, a heterogeneous graph knowledge base for AD, constructed using external data sources and describing various medical entities (e.g., chemicals, genes). Other graph-based efforts model knowledge around AD for tasks such as drug repurposing (Daluwatumulle et al., 2022; Nian et al., 2022; Hsieh et al., 2023), gene identification (Binder et al., 2022), or as general knowledge repositories (Sügis et al., 2019). Another paradigm for knowledge discovery is the open information extraction (OIE) setup (Etzioni et al., 2008; Mausam et al., 2012), which faces challenges such as data consistency, performance evaluation, and semantic drift (Zhou et al., 2022). Research efforts (Nebot and Berlanga, 2011; Movshovitz-Attias and Cohen, 2012; Nebot and Berlanga, 2014; de Silva et al., 2017; Wang et al., 2018) aim to address these problems and extract knowledge with little or no supervision. Advances in literature-based discovery (Gopalakrishnan et al., 2019; Thilakaratne et al., 2019) try to identify novel medical entity relations using graph-based (Kilicoglu et al., 2020; Nicholson and Greene, 2020), machine learning (Zhao et al., 2021; Lardos et al., 2022), and co-occurrence methods (Kuusisto et al., 2020; Millikin et al., 2023). Tian et al. (2024) stress the potential of large language models (LLMs) to summarize, simplify, and synthesize medical evidence (Peng et al., 2023; Tang et al., 2023; Shaib et al., 2023), suggesting that LLMs may have encoded biomedical knowledge (Singhal et al., 2023). To exploit this potential, we explore constructing LM representations for knowledge discovery. To the best of our knowledge, no systematic approach assembles knowledge about RS. Unlike previous work, we introduce a, in principle, disease-agnostic framework, to acquire knowledge about RS and AD starting from raw text.
@@ Appendix A Data Pipeline: Additional Information
To facilitate effective abstract retrieval, we implement an iterative approach to circumvent the API’s limitation of retrieving only 10,000 article IDs per query. This iterative process enables us to access a comprehensive set of PubMed IDs (PMIDs) related to the query. The detailed list of the 82 semantic types of the MetaMapLite-based pipeline is presented in Table 4. In addition to the MetaMapLite-based pipeline, we propose a second pipeline that is based on ScispaCy (Apache License 2.0) (Fig. 5). The difference lies in the selection of entity extractors and linkers that map the extracted entities to knowledge schemes. Unlike MetaMapLite, which adopts an integrated approach where mention extraction and linking are performed simultaneously in a single step and focuses on UMLS mapping, allowing for more precise and targeted extraction of entities, ScispaCy serves a broader range of Natural Language Processing (NLP) tasks. After the retrieval of the abstracts that is described in subsection 2.1, the following steps are executed: Knowledge schema and linker generation, Mention extraction, Entity linking and Sampling of linked identifiers.
@@
Entity linking. This process enhances the semantic understanding of the extracted entities, facilitates standardization, which is a key challenge in the biomedical field (Bettencourt-Silva et al., 2012; Theodoropoulos et al., 2023), and promotes interoperability with external resources by associating the entities with specific concepts in supported knowledge schemes. Each entity is subjected to a linking process where we attempt to map it to concepts within supported knowledge bases or vocabularies. If a match is found, the entity is assigned a unique identifier, referred to as a CUI, corresponding to the specific concept in the knowledge schema. As entities may be linked to multiple knowledge sources, we merge the extracted CUIs obtained from the different linkers. This consolidation process ensures that each entity is associated with a comprehensive set of identifiers, encompassing diverse perspectives and representations across various knowledge schemes.
@@
$sam\_sent\_2\leftarrow sample(\mathcal{IP},n/2)$ return $sam\_sent\_1$, $sam\_sent\_2$
@@ Appendix D Localized Context Vector
Data Type1 Binary setup Multi-class setup
Micro Evaluation Macro Evaluation
BiomedBERT2 BioLinkBERT2 BioGPT BiomedBERT2 BioLinkBERT2 BioGPT BiomedBERT2 BioLinkBERT2 BioGPT
ReDReS A 90.72/89.95 91.15/89.9 90.15/88.83 74.49/73.91 74.06/75.13 72.22/72.65 74.52/74.5 74.11/75.09 71.8/72.95
D 90.55/89.29 90.35/90.12 89.76/88.08 73.61/73.85 74.12/75.19 71.42/73.25 73.71/74.54 74.12/75.54 72.48/73.63
E 89.57/89.39 89.98/90.28 89.49/87.19 73.73/75.67 73.22/ 70.17/70.98 73.95/74.9 73.38/ 69.97/70.69
F 90.48/89.09 90.72/89.84 89.68/88.35 72.86/74.18 74.22/ 71.36/73.57 72.62/72.82 74.13/ 72/73.82
G 90.78/89.32 90.57/89.57 89.93/88.41 74.33/73.35 74/ 72.68/73.14 74.57/73.78 74.32/ 72.48/73.43
O 90.9/89.94 90.75/90.12 89.51/87.82 73.99/74.51 74/ 67.25/68.67 73.83/74.71 73.69/ 67.43/67.75
ReDAD A 88.31/90.15 88.73/89.18 87.49/89.19 77.64/77.07 78.4/76.38 76.04/75.96 77.34/77.21 77.68/76.66 76.45/75.62
D 87.33/88.99 88.74/89.72 87.56/89.63 78.28/76.73 78.27/77.03 76.49/76.02 78.26/76.8 77.89/77.18 76.25/76.04
E 88.03/88.63 88.43/90.48 87.71/87.86 77.83/77.45 78.96/78.02 75.35/73.65 77.75/77.34 78/78.1 75.36/73.28
F 87.71/89.54 88.62/89.86 87.63/88.94 77.59/76.5 77.43/77.25 76.34/75.93 77.35/76.33 77.48/77.45 75.47/76.1
G 88.17/90.06 88.72/90.38 87.55/88.85 77.83/77.64 77.76/76.99 76.35/75.71 77.5/77.69 77.33/76.41 76.61/75.73
O 88.27/90.59 88.58/89.91 86.19/87.74 77.08/79.02 76.89/74.85 /71.32 76.78/78.95 76.62/74.93 70.57/70.75
  • 1
    Type of Relation Representation.
  • 2
    Base version.
Table 5: LaMReDA Results (%) in binary and multi-class setup utilizing BiomedBERT-base, BioLinkBERT-base, and BioGPT as backbone language model: Each cell shows the average F1-score from 10 runs (original test set) and from 5-fold cross-validation setup.
Appendix E Experiments with BioLinkBERT and BioGPT
Assessing the impact of different language models on the performance of the LaMReDA and LaMEL models, we conduct experiments using additional models. Specifically, we evaluate the effect of BioLinkBERT (Yasunaga et al., 2022) (base model) and BioGPT (Luo et al., 2022b), both pre-trained on PubMed abstracts. BioLinkBERT is selected to determine whether leveraging the links between documents in the pre-training corpus benefits the relation extraction task of our study. This model uses citations between PubMed articles to model document connections, which may enhance its ability to capture semantic relations. BioGPT, as the GPT-based (Radford et al., 2018, 2019) transformer decoder (Vaswani et al., 2017) representative, is incorporated to explore the generative model’s capabilities in capturing semantic relations between entities. Yasunaga et al. (2022) state that the pre-training data of BioLinkBERT and BioMedBERT are identical  [^14] https://pubmed.ncbi.nlm.nih.gov. Papers published before Feb. 2020., except that they also use the citations between PubMed articles. Luo et al. (2022b) clarify that BioGPT’s pre-training includes PubMed abstracts updated before 2021. Hence, all language models in our study are pre-trained on identical or nearly identical datasets, making the comparison between them reliable.
We conduct experiments using a variety of representations, specifically testing the $R_{A}$, $R_{D}$, $R_{E}$, $R_{F}$, $R_{G}$, and $R_{O}$ (Eq. 9, 12, 13, 14, 15, 23) relation representations of LaMReDA, as well as the $E_{A}$, $E_{D}$, $E_{E}$, and $E_{H}$ (Eq. 1, 4, 5, 8) entity representations of LaMEL. The experimental setup follows the same procedure outlined in subsection 3.3, and the experiments are performed on both the original split (Tab.1) and in a 5-fold cross-validation setting.
The results in Tables 5 and 6 demonstrate the F1-scores of the LaMReDA and LaMEL models on the ReDReS and ReDAD datasets, using different language models. The experiments show that both LaMReDA and LaMEL perform consistently well across various relations ($R_{A}$, $R_{D}$, $R_{E}$, $R_{F}$, $R_{G}$, $R_{O}$) and entities ($E_{A}$, $E_{D}$, $E_{E}$, $E_{H}$) representations, utilizing the different language models. This indicates the models’ ability to learn meaningful representations for the semantic relation detection task, independent of the initial token selection.
Type1 ReDReS ReDAD
BiomedBERT2 BioLinkBERT2 BioGPT BiomedBERT2 BioLinkBERT2 BioGPT
A 90.25/89.43 90.68/89.02 89.43/ 86.73/88.75 87.71/ 86.14/
D 90.47/88.9 90.35/89.3 88.52/ 86.29/88.88 87.35/ 85.28/
E 90.61/89.1 90.25/89.42 89.48/ 86.03/88.96 87.05/ 87.42/
H 89.68/89.24 90.73/89.48 89.21/ 87.29/89.91 87.56/ 85.55/
  • 1
    Type of Relation Representation.
  • 2
    Base version.
Table 6: LaMEL Results (%) in binary setup utilizing BiomedBERT-base, BioLinkBERT-base, and BioGPT as backbone language model: Each cell shows the average F1-score from 10 runs (original test set) and from 5-fold cross-validation setup.
An intra-lm comparison reveals that the use of relation representation $R_{O}$ results in lower performance across both datasets when BioGPT is employed, particularly in the multi-class setup (Tab. 5). This suggests that the context vector extracted from the generative model is less informative for this task. Conversely, when the context vector is derived from BiomedBERT and BioLinkBERT, the models show strong performance, showing the best results on ReDAD in the 5-fold cross-validation setup (Tab. 5).
The inter-language model comparison indicates that BiomedBERT and BioLinkBERT perform similarly, suggesting that leveraging document links during pre-training, as BioLinkBERT does, does not offer a significant advantage for our task. In contrast, the use of BioGPT leads to lower performance, particularly in the multi-class setup, emphasizing the superiority of encoder-based language models such as BiomedBERT and BioLinkBERT over decoder-based BioGPT for the relation detection task. Despite BioGPT’s larger size, with 347 million parameters (Luo et al., 2022b) compared with the 110 million parameters of BiomedBERT and BioLinkBERT (Vaswani et al., 2017; Yasunaga et al., 2022), the generative approach appears less effective for this particular semantic relation task, further underscoring the advantages of encoder-focused models in this context.
Appendix F Baseline Performance: Lower Bound
@@
Appendix G Error Analysis
Whereas the macro and micro evaluations in the multi-class setup indicate strong overall performance across different relation types, we conduct an error analysis to provide additional qualitative insights. Using the best-performing models on the ReDReS and ReDAD datasets based on the macro evaluation (Tab. 3), we extract predictions from the 5-fold cross-validation setup. Confusion matrices for each class are generated to present a comprehensive view of the misclassified instances.
Figure 8: ReDReS Confusion Matrices for each relation class using the best-performing model with BiomedBERT-base as the backbone language model.
Figure 9: ReDReS Confusion Matrices for each relation class using the best-performing model with BiomedBERT-large as the backbone language model.
Figure 10: ReDAD Confusion Matrices for each relation class using the best-performing model with BiomedBERT-base as the backbone language model.
Figure 11: ReDAD Confusion Matrices for each relation class using the best-performing model with BiomedBERT-large as the backbone language model.
We analyze the false negatives for each relation class to gain insights from the misclassification rates. In the ReDReS dataset, the majority of misclassified instances of the positive relation (324 and 331 instances using BiomedBERT-base and BiomedBERT-large, respectively) are labeled as complex relations (Tab. 7). This is intuitively reasonable, as the model recognizes that the entities are semantically related but fails to predict the correct relation type. Conversely, complex relation instances are often mislabeled as no relation (Tab. 8), highlighting the challenge of detecting nuanced semantic relations that require deeper reasoning. For the no relation class (Tab. 10), most false negatives (364 and 303 instances using BiomedBERT-base and BiomedBERT-large, respectively) are also labeled as complex relations, indicating a model bias toward predicting associations between co-occurring entities. Similar patterns are observed in the ReDAD dataset regarding the positive, complex, and no relation classes. The under-represented negative class has the most misclassified cases labeled as complex or positive relations in ReDReS and ReDAD, respectively, suggesting that while the model detects the semantic connection, it tends to predict one of the two more represented classes in the datasets.
Data BiomedBERT-base BiomedBERT-large
Complex Negative No Relation Complex Negative No Relation
ReDReS 331 14 41 324 11 42
ReDAD 424 27 40 425 21 18
Table 7: False Negative analysis for the positive relation type.
Data BiomedBERT-base BiomedBERT-large
Positive Negative No Relation Positive Negative No Relation
ReDReS 221 26 307 188 8 249
ReDAD 335 19 338 274 6 297
Table 8: False Negative analysis for the complex relation type.
Data BiomedBERT-base BiomedBERT-large
Positive Complex No Relation Positive Complex No Relation
ReDReS 25 28 1 15 34 1
ReDAD 44 23 2 44 25 1
Table 9: False Negative analysis for the negative relation type.
Data BiomedBERT-base BiomedBERT-large
Positive Complex Negative Positive Complex Negative
ReDReS 28 303 2 41 364 0
ReDAD 52 653 1 50 607 1
Table 10: False Negative analysis for the no relation relation type.
Appendix H Distantly Supervised Datasets
@@
  • LaMReDA (BiomedBERT large) with the relation representation $R_{A}$ (Eq. 9) for the RS corpus.
  • LaMReDA (BiomedBERT large) with the relation representation $R_{J}$ (Eq. 18) for the AD corpus.
@@
  • LaMReDA (BiomedBERT large) with the relation representation $R_{F}$ (Eq. 14) for the RS corpus.
  • LaMReDA (BiomedBERT large) with the relation representation $R_{F}$ (Eq. 14) for the AD corpus.
@@
Weakly Supervised Setup. The task formulation remains the same as described in section 3 of the paper. The train sets of ReDReS and ReDAD are replaced by DiSReDReS and DiSReDAD, respectively. The development and test sets remain the same (Tab. 1). To provide a benchmark, we train the LaMReDA (BiomedBERT base) with the relation representation $R_{A}$ (Eq. 9) for 10 epochs utilizing the ADAM optimizer with learning rate 10-5. The batch is set to 32. The experiments are repeated 10 times with different seeds and the best scores are retained based on the performance on the development set.
@@
In the supervised setup (Tab. 1), LaMReDA (BiomedBERT base) with the $R_{A}$ representation achieves 90.72% and 88.31% F1-score in the binary setup on ReDReS and ReDAD, respectively (Tab. 3). Multi-class macro F1-scores range from 74.52% (micro: 74.49%) to 77.34% (micro: 77.64%) for ReDReS and ReDAD, accordingly (Tab. 3). Table 5 presents the benchmark performance in the binary and multi-class setup for both datasets. Notably, the performance is improved in the weakly supervised setup, indicating the robustness of LaMReDA, when trained with noisy data, and highlighting the quality of the silver labels of DiSReDReS and DiSReDAD.
Appendix I Probing: Additional Experiments
@@
Figure 12: ReDReS Probing (Multi-class setup, Macro evaluation) (BiomedBERT base): Examines LaMReDA/LaMReDM relation representations ($R_{D}$, $R_{O}$, $R_{P}$) and attention scores from each layer and explores average attention scores of tokens corresponding to each entity towards the other entity across attention heads. Top boundary: best LaMReDA and LaMReDM performance (Tab. 3). Second boundary: classifier with average attention scores across all layers as input.
We use the same experimental setup as described in subsection 3.3 and the experiments are conducted in the 5-fold cross-validation setting. To provide an inclusive probing analysis on ReDReS, we incorporate additional probing results in this section. Fig. 8 and 9 present the experiments in the multi-class setup using BiomedBERT base. Additionally, aiming to explore the probing capabilities of BiomedBERT large, we include the results of further experiments in Fig. 10, 11, and 12. These experiments investigate the model’s performance in detecting semantic relations, comparing the representations and attention mechanisms at different layers and heads to understand how well the larger LM can discern complex relationships in the biomedical text.
@@
Figure 13: ReDReS Probing (Multi-class setup, Micro evaluation) (BiomedBERT base): Examines LaMReDA/LaMReDM relation representations ($R_{D}$, $R_{O}$, $R_{P}$) and attention scores from each layer and explores average attention scores of tokens corresponding to each entity towards the other entity across attention heads. Top boundary: best LaMReDA and LaMReDM performance (Tab. 3). Second boundary: classifier with average attention scores across all layers as input.
@@
Figure 14: ReDReS Probing (Binary setup) (BiomedBERT large): Examines LaMReDA/LaMReDM relation representations ($R_{D}$, $R_{O}$, $R_{P}$) and attention scores from each layer and explores average attention scores of tokens corresponding to each entity towards the other entity across attention heads. Top boundary: best LaMReDA and LaMReDM performance (Tab. 3). Second boundary: classifier with average attention scores across all layers as input.
@@
Figure 15: ReDReS Probing (Multi-class setup, Macro evaluation) (BiomedBERT large): Examines LaMReDA/LaMReDM relation representations ($R_{D}$, $R_{O}$, $R_{P}$) and attention scores from each layer and explores average attention scores of tokens corresponding to each entity towards the other entity across attention heads. Top boundary: best LaMReDA and LaMReDM performance (Tab. 3). Second boundary: classifier with average attention scores across all layers as input.
@@
Figure 16: ReDReS Probing (Multi-class setup, Micro evaluation) (BiomedBERT large): Examines LaMReDA/LaMReDM relation representations ($R_{D}$, $R_{O}$, $R_{P}$) and attention scores from each layer and explores average attention scores of tokens corresponding to each entity towards the other entity across attention heads. Top boundary: best LaMReDA and LaMReDM performance (Tab. 3). Second boundary: classifier with average attention scores across all layers as input.
Appendix J Detailed Annotation Guidelines